OpenStudy (anonymous):

4/5ths divided by 16/15ths

OpenStudy (twopointinfinity):

Keep, Change, Flip. \[\frac{ 4 }{ 5 } * \frac{ 15 }{ 16 }\] Now all you have to do is multiply straight across

OpenStudy (twopointinfinity):

\[\frac{ (4*15) }{ (5*16) }\]